OpenStudy (anonymous):

use the function f(x)= 3x^2 - 4x +1. find the derivative f(x)

OpenStudy (aravindg):

ok Just find derivatives of each term

OpenStudy (aravindg):

derivative of 3x^2= .....

OpenStudy (aravindg):

d/dx(-4x)=....

OpenStudy (aravindg):

d/dx(1)=......

OpenStudy (aravindg):

Can you find them ?

OpenStudy (anonymous):

how do you find d

OpenStudy (aravindg):

d for ?

OpenStudy (anonymous):

d/dx(1)

OpenStudy (aravindg):

1 is a constant .Derivative of constant=0

OpenStudy (aravindg):

What about other 2 terms ?Did you get them ?

OpenStudy (anonymous):

well i got 6x as my answer after setting it up as 3(x+h)^2 -4x +1 -(3x^2-4x+1)

OpenStudy (aravindg):

6x is right for first term ..But you missed 1 term

OpenStudy (anonymous):

4(x+h)?

OpenStudy (aravindg):

Well Are you required to find answer using this method only ?

OpenStudy (aravindg):

Anyways its -4(x+h) instead of -4x

OpenStudy (anonymous):

so the answer is 6x -4 instead of just 6x

OpenStudy (aravindg):

yes :)

OpenStudy (aravindg):

Great work ! Here is a medal for doing it yourself :)

OpenStudy (anonymous):

i have one other question, do you mind trying

OpenStudy (aravindg):

Yep .But post it separately :)

OpenStudy (aravindg):

Btw Welcome to Openstudy !